Kevin is twice as old as gabriela. 12years ago,kevin was 6times as old as gabriela.how old is gabriela now?
Answer by Theo(13342)

kevin is twice as old as gabriela.
12 years ago, kevin was 6 times as old as gabriela.
how old is gabriela today?

let k = kevin's age today.
let b = gabriela's age today.

k = 2g

this means kevins age is two time gabriela's age today.

12 years ago, kevin was 6 times as old as gabriela was 12 years ago.

kevin's age 12 years ago was k - 12.
gabriela's age 12 hears ago was g - 12.

(k - 12) = 6 * (g - 12)

this means that kevin's age 12 years ago was 6 times gabriela's age 12 years ago.

so, you have 2 equations.

k = 2g
(k - 12) = 6 * (g - 12)

these 2 equations have to be solved simultaneously because the same answer has to apply to both equations and not just one of them.

since k = 2g from the first equation, then replace k in the second equation with 2g.

you get (2g - 12) = 6 * (g - 12)

now this equation can be solved for g.

simplify this equation to get 2g - 12 = 6g - 72

subtract 2g from both sides of the equation and add 72 to both sides of the equation to get 60 = 4g.

solve for g to get g = 60/4 = 15.

if g = 15, then k = 2g = 30.

your answer is that kevin is 30 years old today and gabriela is 15 years old today.

12 years ago, kevin would have been 30 - 12 = 18.
12 years ago, gabriela would have been 15 - 12 = 3.

12 years ago kevin would have been 18 / 3 = 6 times as old as gabriela was 12 years ago.

this confirms the answer is correct because all the problem statements have been satisfied.